Q. 1 If you use 3 mol of KClO3 then how many mole of O2 will produce according to the balance equation 2KClO3 ------à 2KCl +   3 O2                             

Q. 2    How many moles of water H2O are in 75.0 g water?                

Q. 3 One mole of any substance has how many particles?                  

Q. 4   One mole of Helium gas weighs                                                      

Q. 5   Calculate the molar mass of   Na3PO4          

                         

         

( Q 6 -8) Iron oxide (Fe2 O3 ) reacts with CO to form Iron metal (Fe) and CO2

      Fe2O3 + 3 CO ---à    2 Fe    +    3 CO2

If 5.0 g of   Fe2O3 is reacted with 5.0 g CO gas

Q.6 What will be the Limiting Reactant ?                                                   

Q.7 What mass of Fe will produce?                                                            

Q.8     What is the percent of yield of Fe if actual amount is 2.95 gm.     
Answer #1

1)

From balanced equation,

Mol of O2 produced = (3/2)*mol of KClO3 reacted

= (3/2)*3 mol

= 4.5 mol

Answer: 4.5 mol

2)

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass(H2O)= 75.0 g

use:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(75 g)/(18.02 g/mol)

= 4.163 mol

Answer: 4.16 mol

Only 1 question at a time please